OpenStudy (anonymous):

How many and of what type are the solutions to x2 + 7x + 10 = 0? No real solutions Two identical rational solutions Two different rational solutions Two irrational solutions

OpenStudy (anonymous):

the answer is D

OpenStudy (anonymous):

did you find the solution???

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

can you show the solution??

OpenStudy (anonymous):

the answer is not D

OpenStudy (anonymous):

oh sorry i mean C

OpenStudy (anonymous):

x² + 7x + 10 = 0 a= 1, b=7 ,c=10 D=b² - 4ac D= 49-40 D= 9 >0 Two different rational solutions( option c)

OpenStudy (anonymous):

yes option C is correct